Tom is n years old. In 7 years he will be 21 years old. How old is Tom ?

Answer by lenny460(1073)

Tom is n years old. In 7 years he will be 21 years old. How old is Tom ?
n + 7 = 21
Subtract 7 from each side of the equation.
n + 7 - 7 = 21 - 7
Combine like terms:
7 - 7 = 0
21 - 7 = 14

Answer:
Tom is 14 years old.






Lennox Obuong
Algebra Tutor
Nairobi, Kenya
Email: lennoxobuong@yahoo.com